trains a and b are traveling in the sam direction on paralle tracks train a is traveling at 100 miles per hour and train b is traveling at 120 miles perhour. Train a passes at station at 8:20 a.m. If train b passes the same station at 8.32 a.m at what time will train b catch up to train a?
Answer by stanbon(75887)

Trains a and b are traveling in the same direction on parallel tracks.
Train a is traveling at 100 miles per hour and train b is traveling at 120 miles perhour.
-----------------------------
Train a passes a station at 8:20 a.m. If train b passes the same station at 8.32 a.m at what time will train b catch up to train a?
-------------------------------
A train:
rate = 100 mph ; distance = x miles ; time = x/100 hrs
---------------
B train:
rate = 120 mph ; distance = x miles ; time = x/120 hrs
---------------
Equation:
A time - B time = 12/60 hr
---
x/100 - x/120 = 1/5
---
Multiply thru by 1200 to get:
12x - 10x = 240
2x = 240
x = 120 miles
======================
Cheers,
Stan H.
======================